OpenStudy (aaronq):

When functions are said to be orthogonal, is that just a fancy way of saying that they're perpendicular?

OpenStudy (anonymous):

I think orthogonal is more in depth as it is the property of vectors, functions etc. Perpendicular only applies to geometry(?).

OpenStudy (anonymous):

But if you apply perpendicular to geometry, I guess it's quite fancy saying orthogonal. :D

OpenStudy (aaronq):

okay cool! thanks!

OpenStudy (anonymous):

Two vectors x,y are orthogonal if the inner product space is 0. \[\left \langle x,y \right \rangle=x*y=0\] Perpendicular is a special case of orthogonality when \[x,y\neq \varnothing\]

OpenStudy (aaronq):

\(<x,y>\) means dot products?

OpenStudy (aaronq):

product*

OpenStudy (aaronq):

thanks for the explanation !

OpenStudy (anonymous):

In the context of euclidian space, yes.

OpenStudy (anonymous):

No problem! :)